OpenStudy (anonymous):

plz help i really need this credit Consider the expression (5x^3-8x^4)/square root of -8x+5 A.(x+5)[(-8x+5)^(1/2)] B.(x^3)[(-8x+5)^(1/2)] C. (x^3)[8x-5)^(1/2)] D. (x^3)(-8x+5)

OpenStudy (green_1):

hi i'm going to get some friends and experts to see if we can help you

OpenStudy (anonymous):

ok thank you

OpenStudy (green_1):

try using an equation solver on the internet because I cant open up the graph from my computer

OpenStudy (anonymous):

ok i'll try mathway

OpenStudy (anonymous):

thanks for trying i really appreciate it

OpenStudy (anonymous):

\[\frac{ 5x^3-8x^4 }{ \sqrt{-8x+5} }\\=\frac{ 5x^3-8x^4 }{ \sqrt{-8x+5} }\times \frac{ { \sqrt{-8x+5} } }{ { \sqrt{-8x+5} } } \]\[=\frac{ (5x^3-8x^4)(\sqrt{-8x+5} }{ -8x+5}\]\[=x^3\sqrt{-8x+5}\\=x^3(-8x+5)^\frac{ 1 }{ 2 }\\ \therefore\ option\ B \ is\ correct.\]